# Next-order asymptotics for the volume of Schatten balls

### v1

@@ Abstract.
The volume of the unit balls of self-adjoint finite-dimensional Schatten $p$-classes of $n\times n$-matrices, $1\leq p\leq\infty$, is only known exactly for $p=2$ and $p=\infty$. We give an asymptotic expansion of the logarithmic volume to order $o(n)$ for general $p\geq\frac{3}{2}$. The proof rests on asymptotics for the partition function of $\beta$-ensembles due to Leblé and Serfaty [Invent. Math. 210(3):645–757, 2017]. Independently, the case $p\geq 2$ was obtained by Dworaczek Guera, Memin and Pain [arXiv:2511.05386]. In the complex case the asymptotic expansion is continued to order $O(1)$ for all $p\geq 1$.
@@ 1. Introduction
The Schatten or Schatten-von Neumann class consists of all compact operators between two Hilbert spaces whose sequence of singular values belongs to the Lebesgue sequence space $\ell_{p}$. It can be seen a non-commutative version of $\ell_{p}$ and is named after Schatten and von Neumann Sch46 ; Sch50 ; SVN46 . We refer to Pie07 ; PX03 for an overview of their role in functional analysis and Banach space geometry.
In this article, we study the geometry of finite-dimensional Schatten classes and in particular their unit balls. These have been investigated as test cases for conjectures in asymptotic geometric analysis DFG+23 ; GMP25 ; GP07 ; KMP98 ; RV20 , in the context of low-rank matrix recovery and information-based complexity CR12 ; CK15 ; HPV17 ; HPV21 ; PS22 or in quantum information theory ASW11 ; Wil13 . Let $\mathbb{F}_{1}=\mathbb{R}$, $\mathbb{F}_{2}=\mathbb{C}$ and $\mathbb{F}_{4}=\mathbb{H}$ and denote the singular values of a matrix $A\in\mathbb{F}_{\beta}^{n\times n}$, where $\beta\in\{1,2,4\}$, by
@@
For self-adjoint matrices the Schatten $p$-norm $\|A\|_{p}$ becomes a $p$-norm of eigenvalues. In case of non-self adjoint matrices one can also study the rectangular case, see e.g. JKP24 .
@@
where $\Gamma(x)=\int_{0}^{\infty}t^{x-1}e^{-t}\mathrm{d}t$ denotes the Gamma function. For $p=\infty$ it follows from Saint Raymond’s work SR84 and Selberg’s integral formula And91 ; Sel44 that
@@
$\mathrm{vol}(\mathbb{B}_{\infty,\beta}^{n})=2^{d_{n}}(2\pi)^{\beta n(n-1)/4}\prod_{j=0}^{n-1}\frac{\Gamma(1+j\frac{\beta}{2})^{2}\Gamma((j+1)\frac{\beta}{2})}{\Gamma(2+(n+j-1)\frac{\beta}{2})\Gamma(j\frac{\beta}{2})}.$ (3)
For general $1\leq p\leq\infty$, exact volumes are unknown and instead their asymptotic behavior has been studied. Saint Raymond SR84 derived $A(p)>0$ such that, for $\beta\in\{1,2\}$,
@@
and determined $A(2)=e^{-1/4}$ and $A(\infty)=\frac{1}{2}$. Guédon and Paouris GP07 showed that $\mathrm{vol}(B_{p,4}^{n})^{1/4n^{2}}$ and $\mathrm{vol}(\mathbb{B}_{p,\beta}^{n})^{1/d_{n}},\beta\in\{1,2,4\}$ are of the same order $n^{-1/2-1/p}$. Kabluchko, Prochno and Thäle KPT20a determined
@@
and moreover showed in KPT20b that, for $\beta\in\{1,2,4\}$,
@@
We remark that one way to see that $A(p)$ is decreasing in $p$ is to observe that the same is true for the normalized volume $\mathrm{vol}(n^{1/p}\mathbb{B}_{p,\beta}^{n})$. Recently, Dadoun, Fradelizi, Guédon and Zitt (DFG+23, , Remark 4.8) showed (4) for $\beta=4$. Thus, the asymptotics in (4) and (6) hold in all cases $\beta\in\{1,2,4\}$ and $1\leq p\leq\infty$.
In case of $p=\infty$, the quantity $A(\infty)=\frac{1}{2}$ can be expressed in terms of the minimal logarithmic energy of probability measures supported in [1,1][-1,1] attained by the arcsine distribution, see ST97 . We refer to BDF+24 and Bra24 for related asymptotics for discrete minimal logarithmic energy.
@@
Let $\beta\in\{1,2,4\}$. If $p\geq\frac{3}{2}$, then, as $n\to\infty$,
@@
Moreover, if $\beta=2$ and $p\geq 1$, then, for some constant $M_{p}\in\mathbb{R}$,
$\ln\mathrm{vol}(\mathbb{B}_{p,2}^{n})=-\Big(\frac{1}{2}+\frac{1}{p}\Big)n^{2}\ln n+\Big(\frac{1}{2}\ln 2\pi+\frac{3}{4}+\ln A(p)\Big)n^{2}-\ln n+M_{p}+o(1).$ (9)
In case of $p=2$ it is known that $\mathrm{Ent}(\mu_{2})=\ln\pi-\frac{1}{2}$ where $\mu_{2}$ is the semi-circle law. Otherwise the entropy $\mathrm{Ent}(\mu_{p})$ of the Ullman distribution appears to be unknown. Independently, the expansion in (1) was recently obtained by Dworaczek Guera, Memin and Pain GMP25 in case of $p\geq 2$.
@@
$Z_{n,p,\beta}=\int_{\mathbb{R}^{n}}\prod_{1\leq i<j\leq n}|x_{i}-x_{j}|^{\beta}e^{-\frac{\beta}{2}n\,v_{p}\sum_{i=1}^{n}|x_{i}|^{p}}\mathrm{d}x,$ (10)
@@
$v_{p}=\frac{\sqrt{\pi}\Gamma(\frac{p}{2})}{\Gamma(\frac{p+1}{2})}=\frac{1}{p\alpha_{p}}.$
The quantity $Z_{n,p,\beta}$ is the partition function of a one-dimensional $\beta$-ensemble with potential $V(x)=v_{p}|x|^{p}$, where eigenvalues/particles are distributed with joint density
$f_{n,p,\beta}(x)=\frac{1}{Z_{n,p,\beta}}\prod_{1\leq i<j\leq n}|x_{i}-x_{j}|^{\beta}e^{-\frac{\beta}{2}n\,v_{p}\sum_{i=1}^{n}|x_{i}|^{p}}\mathrm{d}x,\qquad x\in\mathbb{R}^{n}.$ (11)
The parameter $\beta$ stands for the inverse temperature and may take any positive value, while for the matrix representation we restrict to $\beta\in\{1,2,4\}$. We refer to AGZ10 and For10 for more information on $\beta$-ensembles.
@@
over $\mu\in\mathcal{P}(\mathbb{R})$, where $\mathcal{P}(\mathbb{R})$ denotes the space of probability measures on $\mathbb{R}$. More precisely, if $X=(X_{1},\dots,X_{n})$ is distributed according to the density $f_{n,p,\beta}$, then the random empirical measure $\mu_{n,p}=\frac{1}{n}\sum_{i=1}^{n}\delta_{X_{i}}$ tends to $\mu_{p}$ as $n\to\infty$. Moreover, large deviation principles due to BG97 and HP00 yield
@@
Earlier work DPS95 ; Joh98 also identified the limiting constant but with different techniques. In DFG+23 , the limit (12) is used to prove (6). The large deviation principle for tagged empirical fields due to Leblé and Serfaty LS17 allows to refine (12). The following result forms the core of the proof of (1) in Theorem 1.
@@
Independently, in the recent work GMP25 the authors deduced the statement of Theorem 2 in case of $p\geq 2$ from a central limit theorem for linear eigenvalue statistics. The restriction to $p\geq 2$ results from regularity requirements of the underlying approach, see also DFG+23 . The slightly milder restriction on $p$ in Theorem 2 is due to the fact that the Ullman density $f_{p}$ is Hölder continuous of order $\frac{1}{2}$ if and only if $p\geq\frac{3}{2}$. This is required to apply the results in LS17 .
The term $\mathrm{Ent}(\mu_{p})$ in Theorem 2 arises via a rescaling property of the specific entropy which is part of the rate function in the large deviation principle in LS17 . In case of $\beta=2$, the contribution of $\mathrm{Ent}(\mu_{p})$ vanishes and more precise asymptotics hold. The following is a consequence of (CKM23, , Prop. 1.1) which is based on KM99 .
@@
Let $p\geq 1$. There exists $M_{p}^{\prime}\in\mathbb{R}$ such that, as $n\to\infty$,
$\ln Z_{n,p,2}=-n^{2}I_{p}(\mu_{p})+n\ln n+(\ln 2\pi-1)n+\frac{5}{12}\ln n+M_{p}^{\prime}+o(1).$
Theorem 3 combined with Lemma 8 and Lemma 10 below yields (9) in Theorem 1. In case of $p$ an even integer, the potential $V(x)=v_{p}|x|^{p}$ is an analytic function and results of BG13 yield a full asymptotic expansion of $\ln Z_{n,p,\beta}$, and consequently of $\ln\mathrm{vol}(\mathbb{B}_{p,\beta}^{n})$. In case of $p=2$ the integral $Z_{n,2,\beta}$ is known via a Mehta-Selberg integral. For $p\neq 2$ the coefficients in the full asymptotic expansion are defined in terms of integrals of the Stieltjes transform of $\mu_{p}$, see also (BG24, , Section 7) and Shc13 ; Shc14 .
In the case of non self-adjoint matrices, the volume of the unit ball $B_{p,\beta}^{n}$ may be written in terms of
$\tilde{Z}_{n,p,\beta}=\int_{(\mathbb{R}_{+})^{n}}\prod_{1\leq i<j\leq n}|x_{i}-x_{j}|^{\beta}\prod_{i=1}^{n}x_{i}^{\beta/2-1}e^{-\beta n\,v_{p}\sum_{i=1}^{n}x_{i}^{p/2}}\mathrm{d}x,$
see e.g. (DFG+23, , Sec. 4.2). This corresponds to the partition function of a Laguerre $\beta$-ensemble with potential $V(x)=v_{p}|x|^{p/2}$. The leading order asymptotics for $\tilde{Z}_{n,p,\beta}$ can again be obtained via large deviations, leading to (4). However, the results in LS17 do not apply and it goes beyond the scope of the present work to derive next-order asymptotics for $\tilde{Z}_{n,p,\beta}$ and $B_{p,\beta}^{n}$.
Remark 4.
The unit balls $B_{p,\beta}^{n}$ and $\mathbb{B}_{p,\beta}^{n}$ can also be studied for $0<p<1$. Then (4) and (6) as well as Theorem 3 continue to hold. However, even neglecting regularity assumptions, the proof of Theorem 1 and the underlying large deviation principle in LS17 are limited to $p\geq 1$. This is because, while the microscopic limit point process in the bulk is the universal sine2-process for $p\geq 1$, it becomes dependent on $p$ for $p<1$, see BEY14 ; CKM23 . Since the minimizer of the rate function is the free energy which is uniquely given by the sineβ-process, see EHL21 ; HM25 , the large deviation principle in LS17 cannot hold in its present form if $p<1$.
@@
where $\nu_{p}$ has density proportional to $e^{-|x|^{p}/p}$ and maximizes entropy over $\mu\in\mathcal{P}(\mathbb{R})$ with $\int|x|^{p}\mathrm{d}\mu(x)\leq 1$, see also KP21 . Similarly, the Ullman distribution $\mu_{p}$ maximizes free entropy over $\mu\in\mathcal{P}(\mathbb{R})$ with $\int|x|^{p}\mathrm{d}\mu(x)\leq\alpha_{p}$, see (HP00, , Prop. 5.34). In case of $p=\infty$, one can interpret $\nu_{\infty}$ as the uniform distribution on [1,1][-1,1], which maximizes entropy, and $\mu_{\infty}$ as the arcsine distribution, which maximizes free entropy. Formally, see Lemma 8 below, the case $\beta=0$ corresponds to the commutative case where a Poisson point process appears in the microscopic limit, see also AD14 ; BP15 ; Leb16 .
@@
Theorem 1 can be combined with the stochastic representation in [Cor. 4.3]KPT20b of a random matrix drawn uniformly from a Schatten ball $\mathbb{B}_{p,\beta}^{n}$ to obtain information on critical intersections of Schatten $p$-balls in the spirit of KPT20b and ST25 .

### v2

@@ Abstract.
The volume of the unit balls of self-adjoint finite-dimensional Schatten $p$-classes of $n\times n$-matrices, $1\leq p\leq\infty$, is only known exactly for $p=2$ and $p=\infty$. We give an asymptotic expansion of the logarithmic volume to order $o(n)$ for all $p>1$. The proof rests on asymptotics for the partition function of $\beta$-ensembles due to Leblé and Serfaty [Invent. Math. 210(3):645–757, 2017]. Independently, the case $p\geq 2$ was obtained by Dworaczek Guera, Memin and Pain [arXiv:2511.05386]. In the complex case the asymptotic expansion is continued to order $o(1)$ for all $p\geq 1$.
@@ 1. Introduction
The Schatten or Schatten-von Neumann class consists of all compact operators between two Hilbert spaces whose sequence of singular values belongs to the Lebesgue sequence space $\ell_{p}$. It can be seen a non-commutative version of $\ell_{p}$ and is named after Schatten and von Neumann Schatten (1946, 1950); Schatten and von Neumann (1946). We refer to Pietsch (2007); Pisier and Xu (2003) for an overview of their role in functional analysis and Banach space geometry.
In this article, we study the geometry of finite-dimensional Schatten classes and in particular their unit balls. These have been investigated as test cases for conjectures in asymptotic geometric analysis Dadoun et al. (2023); Dworaczek Guera et al. (2025); Guédon and Paouris (2007); König et al. (1998); Radke and Vritsiou (2020), in the context of low-rank matrix recovery and information-based complexity Candes and Recht (2012); Chávez-Domínguez and Kutzarova (2015); Hinrichs et al. (2017, 2021); Prochno and Strzelecki (2022) or in quantum information theory Aubrun et al. (2011); Wilde (2013). Let $\mathbb{F}_{1}=\mathbb{R}$, $\mathbb{F}_{2}=\mathbb{C}$ and $\mathbb{F}_{4}=\mathbb{H}$ and denote the singular values of a matrix $A\in\mathbb{F}_{\beta}^{n\times n}$, where $\beta\in\{1,2,4\}$, by
@@
For self-adjoint matrices the Schatten $p$-norm $\|A\|_{p}$ becomes a $p$-norm of eigenvalues. In case of non-self adjoint matrices one can also study the rectangular case, see e.g. Juhos et al. (2024).
@@
where $\Gamma(x)=\int_{0}^{\infty}t^{x-1}e^{-t}\mathrm{d}t$ denotes the Gamma function. For $p=\infty$ it follows from Saint Raymond’s work Saint Raymond (1984) and Selberg’s integral formula Anderson (1991); Selberg (1944) (see also Remark 8 below) that
@@
$\mathrm{vol}(\mathbb{B}_{\infty,\beta}^{n})=2^{d_{n}}(2\pi)^{\beta n(n-1)/4}\prod_{j=0}^{n-1}\frac{\Gamma(1+j\frac{\beta}{2})^{2}}{\Gamma(2+(n+j-1)\frac{\beta}{2})}.$ (3)
For general $1\leq p\leq\infty$, exact volumes are unknown and instead their asymptotic behavior has been studied. Saint Raymond Saint Raymond (1984) derived $A(p)>0$ such that, for $\beta\in\{1,2\}$,
@@
and determined $A(2)=e^{-1/4}$ and $A(\infty)=\frac{1}{2}$. Guédon and Paouris Guédon and Paouris (2007) showed that $\mathrm{vol}(B_{p,4}^{n})^{1/4n^{2}}$ and $\mathrm{vol}(\mathbb{B}_{p,\beta}^{n})^{1/d_{n}},\beta\in\{1,2,4\}$ are of the same order $n^{-1/2-1/p}$. Kabluchko, Prochno and Thäle Kabluchko et al. (2020a) determined
@@
and moreover showed in Kabluchko et al. (2020b) that, for $\beta\in\{1,2,4\}$,
@@
We remark that one way to see that $A(p)$ is decreasing in $p$ is to observe that the same is true for the normalized volume $\mathrm{vol}(n^{1/p}\mathbb{B}_{p,\beta}^{n})$. Recently, Dadoun, Fradelizi, Guédon and Zitt (Dadoun et al., 2023, Remark 4.8) showed (4) for $\beta=4$. Thus, the asymptotics in (4) and (6) hold in all cases $\beta\in\{1,2,4\}$ and $1\leq p\leq\infty$.
In case of $p=\infty$, the quantity $A(\infty)=\frac{1}{2}$ can be expressed in terms of the minimal logarithmic energy of probability measures supported in [1,1][-1,1] attained by the arcsine distribution, see Saff and Totik (1997). We refer to Beltrán et al. (2024) and Brauchart (2024) for related asymptotics for discrete minimal logarithmic energy.
@@
The Ullman distribution originated in the theory of orthogonal polynomials on $\mathbb{R}$ with weight $e^{-c|x|^{p}}$ and $A(p)$ is strongly related to the asymptotics of the sequence of leading coefficients, see Saff and Totik (1997) and (Van Assche, 1987, Sec. 4.1).
@@
Let $\beta\in\{1,2,4\}$. If $p>1$, then, as $n\to\infty$,
@@
Moreover, if $\beta=2$ and $p\geq 1$, then
$\ln\mathrm{vol}(\mathbb{B}_{p,2}^{n})=-\Big(\frac{1}{2}+\frac{1}{p}\Big)n^{2}\ln n+\Big(\frac{1}{2}\ln 2\pi+\frac{3}{4}+\ln A(p)\Big)n^{2}-\ln n+M_{p}+o(1),$ (9)
where
$M_{p}=\frac{5}{12}\ln p+\frac{1}{12}\ln 2.$
In case of $p=2$ it is known that $\mathrm{Ent}(\mu_{2})=\ln\pi-\frac{1}{2}$ where $\mu_{2}$ is the semi-circle law. Otherwise the entropy $\mathrm{Ent}(\mu_{p})$ of the Ullman distribution appears to be unknown. Independently, the expansion in (1) was recently obtained by Dworaczek Guera, Memin and Pain Dworaczek Guera et al. (2025) in case of $p\geq 2$.
In addition to Theorem 1 we obtain for $p=\infty$ from the exact formula (3) and asymptotics for the Barnes $G$-function satisfying $G(n+1)=\prod_{k=0}^{n-1}\Gamma(1+k)$, $n\in\mathbb{N}$, that
$\ln\mathrm{vol}(\mathbb{B}_{\infty,2}^{n})=$ $-\frac{1}{2}n^{2}\ln n+\Big(\frac{1}{2}\ln 2\pi+\frac{3}{4}+\ln A(\infty)\Big)n^{2}$
$-\frac{1}{6}\ln n+\frac{1}{12}\ln 2+2\zeta^{\prime}(-1)+o(1),$ (10)
where $\zeta$ is the Riemann zeta function and $\zeta^{\prime}(-1)=\frac{1}{12}-\ln A$ with $A=1.282$... being the Glaisher-Kinkelin constant. The logarithmic term is what we would expect from
$\ln\mathrm{vol}(\mathbb{B}_{p,\beta}^{n})\leq\ln\mathrm{vol}(\mathbb{B}_{\infty,\beta}^{n})\leq\ln\mathrm{vol}(\mathbb{B}_{p,\beta}^{n})+\frac{d_{n}}{p}\ln n,$
upon setting $p=n^{2}(\ln n)^{2}$ such that the difference between the logarithmic volumes vanishes. The discrepancy in the constant terms in (9) and (1) is because the $o(1)$-term in (9) implicitly depends on $p$.
@@
$Z_{n,p,\beta}=\int_{\mathbb{R}^{n}}\prod_{1\leq i<j\leq n}|x_{i}-x_{j}|^{\beta}\prod_{i=1}^{n}e^{-\frac{\beta}{2}n\,v_{p}|x_{i}|^{p}}\mathrm{d}x,$ (11)
@@
$v_{p}=\frac{\sqrt{\pi}\Gamma(\frac{p}{2})}{\Gamma(\frac{p+1}{2})}=\frac{1}{p\alpha_{p}}$
is chosen for normalization purposes. The quantity $Z_{n,p,\beta}$ is the partition function of a one-dimensional $\beta$-ensemble with potential $V(x)=v_{p}|x|^{p}$, where eigenvalues/particles are distributed with joint density
$f_{n,p,\beta}(x)=\frac{1}{Z_{n,p,\beta}}\prod_{1\leq i<j\leq n}|x_{i}-x_{j}|^{\beta}\prod_{i=1}^{n}e^{-\frac{\beta}{2}n\,v_{p}|x_{i}|^{p}},\qquad x\in\mathbb{R}^{n}.$ (12)
The parameter $\beta$ stands for the inverse temperature and may take any positive value, while for the matrix representation we restrict to $\beta\in\{1,2,4\}$. In case of $p=2$ the integral $Z_{n,2,\beta}$ is known exactly via the Mehta-Selberg integral. We refer to Anderson et al. (2010) and Forrester (2010) for more information on $\beta$-ensembles.
@@
over $\mu\in\mathcal{P}(\mathbb{R})$, where $\mathcal{P}(\mathbb{R})$ denotes the space of probability measures on $\mathbb{R}$. More precisely, if $X=(X_{1},\dots,X_{n})$ is distributed according to the density $f_{n,p,\beta}$, then the random empirical measure $\mu_{n,p}=\frac{1}{n}\sum_{i=1}^{n}\delta_{X_{i}}$ tends to $\mu_{p}$ as $n\to\infty$. Moreover, large deviation principles due to Ben Arous and Guionnet (1997) and Hiai and Petz (2000) yield
@@
Earlier work De Monvel et al. (1995); Johansson (1998) also identified the limiting constant but with different techniques. In Dadoun et al. (2023), the limit (13) is used to prove (6). The large deviation principle for the corresponding tagged empirical fields due to Leblé and Serfaty Leblé and Serfaty (2017) allows to refine (13) as follows.
@@
Theorem 2 forms the core of the proof of (1) in Theorem 1. Independently, in the recent work Dworaczek Guera et al. (2025) the authors deduced the statement of Theorem 2 in case of $p\geq 2$ from a central limit theorem for linear eigenvalue statistics. The restriction to $p\geq 2$ results from regularity requirements of the underlying approach, see also Dadoun et al. (2023). The milder restriction to $p>1$ in Theorem 2 is due to the fact that in this case the Ullman density $f_{p}$ is Hölder continuous inside its support which is already sufficient for the approach in Leblé and Serfaty (2017).
The term $\mathrm{Ent}(\mu_{p})$ in Theorem 2 arises via a rescaling property of the specific entropy which is one part of the rate function in the large deviation principle in Leblé and Serfaty (2017). In case of $\beta=2$, the contribution of $\mathrm{Ent}(\mu_{p})$ vanishes and more precise asymptotics hold. The following is a consequence of (Claeys et al., 2023, Prop. 1.1), derived from Kriecherbauer and McLaughlin (1999), and (Claeys et al., 2023, Thm. 1.9).
@@
Let $p\geq 1$. As $n\to\infty$,
$\ln Z_{n,p,2}=-n^{2}I_{p}(\mu_{p})+n\ln n+(\ln 2\pi-1)n+\frac{5}{12}\ln n-\frac{1}{12}\ln\frac{p}{2}+\zeta^{\prime}(-1)+o(1),$
where $\zeta$ is the Riemann zeta function.
Theorem 3 combined with Lemma 7 and Lemma 9 below yields (9) in Theorem 1. In case of $p$ an even integer, the potential $V(x)=v_{p}|x|^{p}$ is an analytic function and (Borot and Guionnet, 2013, Prop. 1.2) yields a full asymptotic expansion of $\ln Z_{n,p,\beta}$, and consequently of $\ln\mathrm{vol}(\mathbb{B}_{p,\beta}^{n})$, with coefficients defined in terms of integrals of the Stieltjes transform of $\mu_{p}$, see also (Borot and Guionnet, 2024, Sec. 7) and Shcherbina (2013, 2014).
In the case of non self-adjoint matrices, for $\beta\in\{1,2,4\}$ and $1\leq p<\infty$, the volume of the unit ball $B_{p,\beta}^{n}$ may be written in terms of
$\tilde{Z}_{n,p,\beta}=\int_{(\mathbb{R}_{+})^{n}}\prod_{1\leq i<j\leq n}|x_{i}-x_{j}|^{\beta}\prod_{i=1}^{n}x_{i}^{\beta/2-1}\prod_{i=1}^{n}e^{-\beta n\,v_{p}x_{i}^{p/2}}\mathrm{d}x,$
see e.g. (Dadoun et al., 2023, Sec. 4.2). This corresponds to the partition function of a Laguerre $\beta$-ensemble with potential $V(x)=v_{p}|x|^{p/2}$. The leading order asymptotics for $\tilde{Z}_{n,p,\beta}$ can again be obtained via large deviations, leading to (4). However, the results in Leblé and Serfaty (2017) do not apply because of a blow-up of the corresponding equilibrium density near the boundary of its support. It goes beyond the scope of the present work to derive next-order asymptotics for $\tilde{Z}_{n,p,\beta}$ and $B_{p,\beta}^{n}$. For completeness, we note that in case of $p=\infty$ it can be derived that (see e.g. Juhos et al. (2024))
$\ln\mathrm{vol}(B_{\infty,\beta}^{n})=$ $-\frac{\beta}{2}n^{2}\ln n+\beta\Big(\frac{1}{2}\ln\frac{2\pi}{\beta}+\frac{3}{4}+\ln A(\infty)\Big)n^{2}$
$+\Big(1-\frac{\beta}{2}\Big)\ln(A(\infty))n+\frac{-3+\frac{\beta}{2}+\frac{2}{\beta}}{12}\ln n+O(1).$
The unit balls $B_{p,\beta}^{n}$ and $\mathbb{B}_{p,\beta}^{n}$ can even be studied for $0<p<1$. Then (4) and (6) as well as Theorem 3 (with a possibly modified constant term) continue to hold. However, the proof of Theorem 1 and the underlying large deviation principle in Leblé and Serfaty (2017) appear to be limited to $p\geq 1$. This is because the rate function in Leblé and Serfaty (2017) is the free energy which is uniquely minimized by the sineβ-process, see Erbar et al. (2021); Huesmann and Müller (2025). While the microscopic limit point process in the bulk of many $\beta$-ensembles is given by the universal sineβ-process, see Bourgade et al. (2014); Mehta (1991); Valkó and Virág (2009), for $\beta=2$ and $p<1$ the density $f_{p}$ develops a singularity at zero near which the limit point process depends on $p$, see Claeys et al. (2023). In case of $\beta\in\{1,4\}$, it is an open problem to extend Theorem 1 to $p=1$ for which
$f_{1}(x)=\frac{1}{\pi}\ln\frac{1+\sqrt{1-x^{2}}}{|x|},\qquad x\in[-1,1].$
The logarithmic singularity at the origin prevents direct application of the results in Leblé and Serfaty (2017) but possibly behaves well enough such that the approach may be adapted to this case.
@@
where $\nu_{p}$ has density proportional to $e^{-|x|^{p}/p}$ and maximizes entropy over $\mu\in\mathcal{P}(\mathbb{R})$ with $\int|x|^{p}\mathrm{d}\mu(x)\leq 1$, see also Kabluchko and Prochno (2021). Similarly, the Ullman distribution $\mu_{p}$ maximizes free entropy over $\mu\in\mathcal{P}(\mathbb{R})$ with $\int|x|^{p}\mathrm{d}\mu(x)\leq\alpha_{p}$, see (Hiai and Petz, 2000, Prop. 5.34). In case of $p=\infty$, one can interpret $\nu_{\infty}$ as the uniform distribution on [1,1][-1,1], which maximizes entropy, and $\mu_{\infty}$ as the arcsine distribution, which maximizes free entropy. Formally, see Lemma 7 below, the case $\beta=0$ corresponds to the commutative case where a Poisson point process appears in the microscopic limit, see also Allez and Dumaz (2014); Benaych-Georges and Péché (2015); Leblé (2016). Possibly, a careful analysis of the behavior of the asymptotics for $\ln Z_{n,p,\beta}$ as $\beta\to 0$ could allow a computation of $\mathrm{Ent}(\mu_{p})$ in terms of $\mathrm{Ent}(\nu_{p})$.
@@
Theorem 1 can be combined with the stochastic representation in (Kabluchko et al., 2020b, Cor. 4.3) of a random matrix drawn uniformly from a Schatten ball $\mathbb{B}_{p,\beta}^{n}$ to obtain information on critical intersections of Schatten $p$-balls in the spirit of Kabluchko et al. (2020b) and Sonnleitner and Thäle (2025).